OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5

OpenStudy (anonymous):

-4x - 2y = 10 -4y = 4 -4x - 5y = -1 -7y = 5 3x + 2y = 6 7x = 12 4x + 2y = 10 8x = 16